OpenStudy (anonymous):

The 64.8-g sample of the compound X2O5 contains 48.0 g of oxygen atoms. What is the molar mass of element X?

OpenStudy (rushwr):

What do u think ?

OpenStudy (anonymous):

i don't know i tried 5 times & continue to get the wrong answer

OpenStudy (rushwr):

Mass of X2 = 64.8g - 48g right?

OpenStudy (rushwr):

So the mass of X2 is 16.8g

OpenStudy (rushwr):

Now we have to find the moles of oxygen

OpenStudy (rushwr):

\[moles = \frac{ mass }{ molar mass} = \frac{ 48 }{ 16 } = 3moles\]

OpenStudy (rushwr):

are u there?

OpenStudy (anonymous):

yes

OpenStudy (rushwr):

\[\frac{ 2 }{ 5 } = \frac{ xmoles }{ 3 }\] \[x moles = 1.2\] Now we know the moles of X2

OpenStudy (anonymous):

okay..

OpenStudy (rushwr):

we know the moles and the mass we just have to find the molar mass by using molar mass = mass/moles

OpenStudy (rushwr):

\[molar mass = \frac{ 16.8 }{ 1.2 } = 14gmol ^{-1}\]

OpenStudy (anonymous):

YES thank you

OpenStudy (rushwr):

Welcome !!!!!!!!! DId u get it ? :)

OpenStudy (anonymous):

yes!